### v1

# Extreme heat is associated with reductions in human activity

Abstract
Extreme heat is a growing threat to both individual livelihoods and broader economies, killing an increasing number of people each year as temperatures increase and limiting productivity—often in the countries that would benefit most from economic growth. Many studies document the link between heat waves and mortality or morbidity, and others explore the economic consequences of them, but few are able to identify the ways in which populations respond to the shock of extreme heat. Toward this end, we investigate the link between human mobility and fluctuations in temperature. Using additive models fitting data in Indonesia, India and Mexico, we show that extreme heat reduces mobility by up to 10% in rural areas and 5% in urban settings. Effects are stronger in poorer areas. Twinning these models with climate projections, we show that without adaptation mobility will fall 1-2% per year on aggregate, with certain seasons and places seeing activity fall by as much as 10%. According to our estimates, rural areas will face the highest relative losses and urban megacities will experience the greatest absolute impacts.
Introduction
Extreme heat threatens lives annually across the world [1, 2]. Various studies establish a relationship between rising temperatures and increases in death rates [3] and hospital admissions [4]. The worst effects of high temperatures occur in populations that are unable to cope with its effects, including the very young and the very old [1, 5]. In addition to these human costs, heat waves bring substantial economic costs, reducing productivity [6] and slowing growth [7]. This problem will likely get worse: heat waves are increasing in both magnitude and frequency [8], with duration and intensity growing since 1950 [9], and increasing at an increasing rate since 1980. Yet literature exploring the behavioral aspects of heat, which underpin these human and economic costs, has been limited in scope and scale. Here we address this gap by joining mass mobility data across 3 developing countries—India, Indonesia and Mexico—with a combined 1.8 billion people across 8 climate zones [10] with data on temperature, precipitation and solar radiation to understand how extreme heat changes activity patterns in both rural and urban contexts.
However, the literature on extreme heat shows that the impact of heat waves is not uniform across all populations. Differences arise due to two primary factors: where a person lives and who the person is. Regarding the latter, age, gender, and specific physical conditions can significantly influence how individuals are affected by extreme heat [11, 12]. As for the former, some cities—or even parts of cities—are more resilient to heat waves than others due to various factors such as proximity to water [13], urban design [14], and other geographic characteristics [15]. Additionally, in areas where people can afford cars, air conditioning, or other amenities, the impact of heat is less significant [16]. Consequently, heat waves tend to disproportionately affect different socioeconomic groups, which is evident when examining changes in their mobility patterns. For instance, areas with large numbers of disadvantaged social groups experience more severe disruptions in mobility due to temperature changes [16].
Data from the US give a window into the relationship between extreme heat and wealth. The association between heat and mortality has fallen in the US since the diffusion of air conditioning, an indication that wealth alleviates the worst effects of climate change [17]. Evidence suggests that quality of life is highest for Americans at average daily temperatures of $18^{\circ}$, and that a given person would pay more to avoid heat than cold [18]. Time use data indicate that individuals exposed to high temperatures at work reduce hours worked and that outdoor leisure is redirected indoors [19]. While this context is considerably different to that of our study here, these results serve as indications for how affluent groups—who can afford to adapt both environments and behaviors—respond to the stresses of heat.
Differences between socio-economic groups are evident not only in their initial reactions to extreme weather events but also in their recovery dynamics. With the availability of granular data capturing human movement, it has been observed that although wealthier groups initially responded more quickly to certain weather events, residents from poorer areas returned to their prior mobility patterns faster [20]. This may be in part because wealthier individuals have the means to reduce their mobility for longer periods. However, more importantly, it suggests that poorer groups face a higher risk of exposure to extreme weather, as they may be compelled to resume activity sooner, despite hazardous conditions.
Yet mortality decreases as the frequency of high temperatures increases, suggesting that populations are able to compensate or adapt to high, albeit frequent, temperatures [21]. Further, the most frequent temperature in an area is often the temperature at which the fewest deaths occur [22]. The possibility for compensatory behavior has led governments and researchers to classify extreme heat alongside hurricanes and cyclones as disruptive events, going so far as to name heat waves as we do storms [23] and develop “heat action plans” [24] to address the problem.
Extreme heat is associated with falling productivity in manufacturing [6] and agriculture [25]; it reduces economic growth in the poor countries that would benefit most from it [7]. These effects are not limited to durable changes in climate: there is evidence that anomalous hot days, rather than long periods of hot temperatures, are enough to curb productivity [26]. A growing body of evidence also indicates that heat waves cause a variety of other social ills, from increasing conflict [27] to decreasing learning amongst pupils in schools without air conditioning [28]. With further implications for the economy, heat waves cause problems for both the built and natural environments, stressing railways and roadways [29], as well as trees and plants [30].
The mechanisms by which heat influences economic behavior are clear: stress from extreme heat causes dehydration and fatigue [31], with particular risk to the elderly. For professions demanding physical labor, heat stress and heat stroke may cause organ damage or death, and can exacerbate existing health problems [24]. In addition to its impacts on physical ability, heat waves also impact cognitive ability and focus [32]. There is evidence that heat reduces hours worked and limits productivity [33]. It primarily affects agricultural labor but it also affects construction and manufacturing [34].
Research, primarily in the US but of particular relevance to this study, finds a connection between heat waves and urban activity. Rising temperatures can alter the types of activities people choose to engage in, the modes of transport they opt for, the distances they are willing to travel, and the duration of their outings [35, 36]. Notably, research indicates that during periods of heightened temperature, people tend to avoid short trips [16] and visit different types of places [37], suggesting that certain trips are not worth the trouble during extreme heat. They also use public transport less [38] and walk less [39], suggesting that comfort is a factor.
Here we attempt to isolate the behavioral foundations of these economic and human consequences by exploring how human mobility responds to changes in temperature. Research into the behavioral implications of extreme heat is growing in importance as heat waves become frequent [40]. Although studies have been able to identify economic consequences of high temperatures, the behavioral foundations of these consequences—how human behavior responds to heat, producing these effects—remains elusive. The same holds for mortality and morbidity: we know that extreme heat kills, but we know less about how individuals or groups of individuals change planned activities to reduce the likelihood of harm.
Operating under the hypothesis that extreme heat curbs mobility, the following paper asks how extreme heat changes mobility. We explore these behavioral foundations of mortality, morbidity and productivity with mass GPS signal data from mobile phones across 3 countries. We focus on India, Indonesia and Mexico—countries with large populations and different climates. Studying countries that are already hot lends importance to our study: in economic simulations, the coldest parts of the world see productivity rise with temperatures while the hottest parts experience the opposite effect [41]. We further disaggregate our analysis according to population and deprivation in order to understand how different groups in these countries respond to and recover from heat waves. Our approach sheds new light on how people manage extreme heat in developing countries and how socio-economic conditions can help or harm these efforts, and it provides the first attempt to model these behaviors across multiple climates and cultures.
Results
We use daily flows between origin and destination geohash5 aggregations, comprising $\sim 236$ million trips and $\sim 1.4$ billion unique GPS signals, gathered from mobile devices in the course of daily usage (for descriptive statistics and illustrations of the network, see Fig. S1 in the supplemental information). For each geohash5 aggregation, we acquire daily temperature, precipitation and solar radiation data from the ERA5 program [42]—data resolved enough that we can observe the temperature gradient within a region, as well as across the country. Joining the two datasets, we are able to make inferences about the effect of temperature on flows to, from and between geohashes.
To begin, we use a binary classification of extreme heat to establish a connection between temperature and flows, borrowing from earlier work the definition of a heat wave as at least 3 days of top decile temperature for a region [43]. We identify periods that meet this criterion in 2019 in India and Indonesia in Fig. 1A. We use spatial interaction networks for Delhi and Jakarta, where each node is a geohash5 and each edge is the average daily trips between those nodes. We calculate these averages for the period of the heat wave and then for the periods immediately preceding and following it, comparing the within and without heat wave values. The results are shown in Fig. 1B and D: on average, edge weights (trips between geohashes) fell by 20% in Delhi and 30% in Jakarta; in the former there are some edges with gains and others with losses while in the latter the drop in activity is uniform. We use cities that did not experience heat waves as a comparison, shown in Fig. 1C, noting that much of the difference in Jakarta relative to Medan, is the result of origin-destination pairs with some trips then having 0 during the heat wave; Delhi saw a broader shift relative to Mumbai. These experiments, however, are only an initial attempt to understand the relationship between temperature and activity, and do not establish a causal link; next we model the relationship formally.
Figure 1: Extreme heat and mobility. A Temperatures in India and Indonesia throughout 2019, highlighting the hottest months of the year for both countries. B and D Change in mean daily visitation in Delhi and Jakarta between hottest week of the year and the weeks preceding and following it; activity levels fell 30% in Jakarta and 20% in Delhi. The results may be more mixed for Delhi because the hottest week of the year came in the middle of a 5-week heat wave which had already begun to influence activity patterns. C The change in distributions for Delhi and Jakarta, along with control cities that did not experience heat waves: while Delhi sees a broad shift relative to Mumbai, the lower average in Jakarta is driven by origin-destination pairs that fall from some trips to 0 during the heat wave.
Modeling the effect of extreme heat on mobility
Because the network tends to be sparse and our concern in this study is “activity”, we can define it as the number of journeys terminating in a given geohash5. In order to understand the effect of incremental changes in temperature, we model the relationship between temperature and activity using generalized additive models (GAMs), which combine the flexibility of non-parametric models with the interpretability of parametric approaches. GAMs use smooth functions that allow coefficients to vary across different points in the function, and they are appropriate for our study in particular because of the documented nonlinear effects of temperature on economic indicators like productivity [44]. (See Methods section for more detail.) We use a directed acyclic graph (DAG) to identify the minimally sufficient adjustment set—the controls that allow us to infer causal links between our independent variable of interest (temperature) with our dependent variable (activity). Shown in Fig. 2A, we identify temporal and spatial factors (identified with blue), like seasonality, deprivation and population, as determinants of activity, along with other aspects of weather—solar radiation and precipitation. Spatial fixed effects at the level of the geohash5 close all geographic confounds and further controls for weather and seasonality close the rest. In alternative specifications, we relax these fixed effects to allow for the exploration of additional factors like population and deprivation (shown in red).
Our first model considers the effect of temperature fluctuations, relative to the local average. Because prior research indicates that frequent temperatures are less lethal [21, 22], we expect that it is these relative changes rather than absolute values that matter. Fig. 1A and B show the relationship between changes in temperature (by standard deviation) and trip counts and durations (measured in minutes), respectively. Our results agree with existing literature: as temperature increases relative to the local average, activity according to these 2 measures decreases. For trip counts, the results are mixed for decreases in temperature, with strong negative effects in Indonesia and weak positive effects in India.
Figure 2: Modeling temperature and activity. A We set up a directed acyclic graph (DAG) to ensure that we close all necessary causal paths to activity; we control for day-of-year, day-of-week, holidays, solar radiation, precipitation, and use geographic fixed effects that necessarily stratify by population and deprivation. B Model results for different countries, showing that the highest activity levels occur at average temperatures for an area, and that high extremes correspond with fewer trips. C Model results using trip duration rather than trip count reveal that the longest trips tend to occur at average temperatures, with extreme high temperatures leading to shorter trips.
We see in Fig. 2B that deviation from the mean temperature—in either direction—in a location is associated with reductions in activity. Showing that relative changes in temperature are better predictors than absolute changes, our results are consistent with the literature that suggests climate preferences are set locally [18]. This also agrees with research that finds that the most frequent temperature in an area is often the least fatal for that area’s population [22]. Exploring other possible changes to activity, Fig. 2C shows the partial effect of temperature anomaly on trip length: while the number of trips fall during extreme high and low temperatures, the lengths of those trips also fall.
We then look at absolute temperatures to understand if levels are also a factor in driving activity. Because prolonged heat waves could elicit changes in behavior relative to shorter ones, we test our model using rolling averages of different lengths—because higher rolling averages indicate stronger or longer fluctuations in temperature. Fig. 3A shows that in India, when the temperature exceeds $33^{\circ}$, the effect of temperature on activity becomes negative; Fig. 3B shows that in Indonesia, this change from positive to negative occurs at $35^{\circ}$. A key observation for India is that during persistent heat waves, captured with longer rolling averages above $42^{\circ}$, the model projects relative adaptation, with the effect size approaching zero. While inconclusive, this suggests that extended periods of high temperatures necessitate a return-to-activity, or that policy and behavior adapt only after multiple days of extreme heat.
Finally, we are interested in the ways different demographic and geographic factors influence activity. This requires us to relax our fixed effects so that we can control for variance along other dimensions: when we control for minimum spatial level of aggregation, spatial attributes cannot vary within that level. In this specification, we use geohash3 fixed effects but add controls for population [45] and deprivation [46] at the level of geohash5. This model accounts for 57% of the variance, compared to 60%, suggesting we reclaim much of the benefit of earlier models while allowing us to explore further relationships. In order to understand the interaction between temperature and deprivation, we explore the space of model predictions across all combinations of temperature and deprivation. These predictions are shown in Fig. 3C and D, for India and Indonesia, respectively.
In India, we recover the same arc as above, where activity partially recovers at the highest extremes. The contours in 3C are flatter for low deprivation and steeper for high deprivation, indicating that much of the worst effects of heat in India are felt by the poorest. Looking at Indonesia in 3D, higher temperatures bring lower activity for all levels of deprivation, but the steeper contours at higher levels of deprivation indicate that the effect is stronger for poor areas than for wealthy ones.
Figure 3: The effect of extreme heat on activity. A Focusing on India, given the propensity for heat waves in India, we measure the duration and intensity of extreme heat by calculating a rolling average, finding that higher temperatures for longer are associated with stronger effects between $35-40^{\circ}$ but not above that, which may indicate compensatory behavior at a certain point. B We fit the same model for Indonesia and find similar results, with an ideal temperature and declining activity when heat surpasses it (noting that Indonesia does not have the same extremes, as it is situated within the tropics). Lasting heat waves, measured in rolling averages, show less impact in Indonesia. C and D Looking at the interaction between deprivation and temperature in India and Indonesia, respectively, we see a curve that follows earlier modeling for all socio-economic strata but it is steepest, indicating the strongest effects, for the poorest strata.
As a robustness check, we use an ARIMAX model to fit the time series of the data for each geohash3 (because the geohash5 level often has incomplete data across the year). This model predicts activity by considering activity at time $t$ using activity at $t-1$ and $t-7$ along with linear and nonlinear terms for temperature, which allows the relationship to curve as temperature increases. This method implicitly controls for seasonality and day-of-week effects by using the previous day and the same day of the previous week to make a prediction, and achieves a good fit, shown for India in Fig. 4A. Important to this study, the quadratic term is negative across most regions, indicating that at a certain point the effect of temperature on activity turns negative. Fig. 4B shows, again for India, the sign and magnitude for each geohash3, which shows mixed results but less of temperature effect in the South, where temperatures are more consistent, and more of an effect in the North and Northwest. Some coefficients are positive with high magnitude, but these tend to be in the mountainous North, which does not have as many hot days, and the tropical South, which has less variability in temperature. We take the average coefficient for $\text{Temperature}^{2}$ across all countries in Fig. 4C. Note that we need to rescale the trip distribution in each country for the effect sizes to be comparable, so the value of each coefficient is difficult to interpret (in Fig. S2, we plot the curves from each ARIMAX formula to show how $\text{Temperature}^{2}$ influences the prediction). We filter out geohashes according to the length of the time series, because very short series could bias the results; generally, as this threshold of minimum observations rises, so too does the effect size.
Finally, we identify the optimal temperature by making predictions across the spectrum for each geohash3 and observing where the predictions begin to fall. According to these predictions, the mean and median optimal temperatures for India are $31.1^{\circ}$ and $31.8^{\circ}$—in line with our estimates from the GAM. (These mirror what we see in S2, which plots the prediction curves for each geohash3 ARIMAX.) Although not a perfect match with the above findings, these results strengthen our assertion that extreme heat can limit activity. We document a noisy relationship between temperature optima across geohashes and the mean temperature of those geohashes ($\rho=0.25$ across all countries, $p<0.05$) in Fig. 4D: as mean temperature increases, the optimal temperature for activity also increases. This provides further evidence that populations adapt to climatic conditions.
Figure 4: Alternative estimations across all countries. A We fit an ARIMAX model to the time series of each geohash3, achieving good fit with a mean absolute percent error of 20%. This model uses both autoregression—recent experience—and exogenous variation, temperature in our case, to predict activity, allowing us to interpret the effect of temperature; B shows the effect of $\text{Temperature}^{2}$ across India, with most areas showing a negative effect. C As a sensitivity analysis, we vary the required length of the time series, because some areas do not have complete data, and find that the effects only become stronger as we increase the minimum number of days in the series. Although a weak correlation of $\rho=0.25$, in D we see a relationship between average temperature and the optimal temperature for activity according to our model, supporting our earlier results.
Projecting into the future
We next turn to models of Earth’s climate [47] to make estimates of how activity will be impacted in future. In particular we use data from CMIP6 simulations, extracting the estimates from the Max Planck Institute because these estimates show the best fit to past trends [48]. We then simply change the temperature-day value in 2019 with temperature-day value in 2050 according to these models. We make these projections knowing that our own study suggests that optimal temperatures are often related to the climate of the area, and thus that the population of a given area may be able to adapt to local temperatures. Nevertheless, our data suggest that optimal temperature is not always average temperature, indicating that rapid deviation from these temperatures as global warming accelerates [9] might still be a shock to communities who have acclimated to past temperatures. Furthermore, our data suggest that poor areas suffer the worst effects, and these areas might be the least able to cope with changes to the local climate going forward. As a modeling exercise, this provides us with a worst-case-scenario, but a plausible one at that.
According to CMIP6 estimates shown in Fig. 5A we can expect 100-150 days per year across Northern states like Rajasthan to exceed $40^{\circ}$ and those same estimates B suggests 25 days per year will exceed $45^{\circ}$. On any given day in June when such temperatures are most likely, not a weekend nor a holiday, our model estimates that a single day at $40^{\circ}$ temperatures to reduce activity in urban areas by 2% and in rural areas by 6%, shown in Fig. 5C. Because cities generate more activity, they experience a small relative change but a large absolute loss. As we see in Fig. 5D, these losses grow to 4% and 8% for urban and rural areas, respectively, during a single day of temperatures at $40^{\circ}$.
CMIP6 estimates mapped in Fig. 6A show that many days per year will exceed $35^{\circ}$ in Sumatra and Borneo. While temperatures of $40^{\circ}$ will be rarer, shown in B they are forecast to happen despite never appearing in our 2019 data. Because Indonesia is in the tropics and does not experience much temperature variability, our model projects strong effects for heat that exists beyond the current distribution. Indonesia is planning to move its capital to the island of Borneo, from Java, by 2045. While this megaproject intends to curb the effects of subsidence, temperatures in this new region, according to CMIP6 data, will experience more hot days than in the existing capital region. Shown in Fig. 6C and D, the consequences of these temperatures, according to our modeling, are similar to the effects we see in India: rural areas with larger relative effects and urban areas with larger aggregate disruptions.
In the supplemental information (Fig. S3), we show the same results for Mexico, which show the same strong effects from extreme heat and mild effects from temperatures that will occur often. An important difference to note is that while the Southern, tropical areas in Mexico will have the highest average temperature, with Fig.LABEL:S3A showing most days over $35^{\circ}$ by 2050, but the desert near the Northern border—where manufacturing serves the US market—will experience more heat waves, highlighted in Fig. LABEL:S3B. Because the literature indicates some adaptation to heat, these heat waves in the North could disrupt activity more than stably higher temperatures in the South.
Figure 5: Implications for the future. A CMIP6 projections for days above $40^{\circ}$ in India by 2050; these indicate that Northern India could experience as many as 150 days per year at or above this temperature. B CMIP6 projects fewer days above $45^{\circ}$ but much of the North will still experience 25 days at this extreme. C and D show the consequences of this heat according to our modeling: during those days between $40-45^{\circ}$, we can expect activity to decrease by $5\%$ in rural areas and $2\%$ in urban ones; when the temperature is higher, the loss of activity could be as much as $7\%$ and $5\%$ in rural and urban areas respectively.
Figure 6: Implications for the future. A CMIP6 projections for days above $35^{\circ}$ in Indonesia by 2050; these indicate that Sumatra and Borneo could experience as many as 125 days per year at or above this temperature. B CMIP6 projects fewer days above $40^{\circ}$ but parts of Sumatra will experience 50-100 days at this level. C and D show the consequences of this heat according to our modeling: during those days between $35-40^{\circ}$, we can expect activity to decrease by $5\%$ in rural areas and $3\%$ in urban ones; when the temperature is higher, the loss of activity could be as much as $8\%$ and $6\%$ in rural and urban areas respectively.
There is an important caveat when we model the entire year: warm weather during periods and in places that were once cold is good for activity, so in India our model expects aggregate mobility to fall only by 1% if we substitute the weather of 2050. Because Indonesia does not have cold weather, rising temperatures there are associated with an decline of 2% across the year. In Mexico, activity will fall as temperatures rise, by 3% on aggregate. This points to changing dynamics: we can expect heat waves to increase in both frequency and magnitude, and thus for associated disruptions to increase, but rising temperatures will have mixed effects. These effects will be spatially and temporally heterogeneous, affecting cities and seasons in different ways.
Discussion
Heat waves are expected to become a regular occurrence in both rural and urban areas over the coming decades. While efforts to mitigate climate change are important, we must also develop adaptive strategies. Understanding the relationship between human behavior and extreme weather, such as heat waves, is essential for this. Our study considerably expands to a growing body of research by examining how extreme heat affects human mobility in three large developing countries. We findings show that heat waves significantly change both urban and rural mobility patterns. By examining three diverse counties across different climate zones, our research allows for comparisons and helps identify broader behavioral trends. Generally, we observed similar patterns in each context: the highest activity levels and longest trips occur at the mean temperatures for each region. However, the specifics vary between these counties. For instance, in India, activity levels were more affected by extreme high temperatures than by lower ones, whereas in Mexico and Indonesia, the reduction in activities was almost symmetric between high and low temperature extremes.
Because none of our modeling techniques are perfect, we triangulate our results with separate approaches—comparing cities before, during and after heats waves, along with additive and autoregressive models. These all point in the same direction: a nonlinear relationship between heat and activity, often with strong negative effects at the extremes. Less certain are our estimates for local differences that suggest adaptation, as average temperature correlates with optimal temperature along these nonlinear curves, but additive and autoregressive models both point to this effect. Because the relationship is noisy, further research could test the whether long-run averages are more predictive of optimal temperature than the short-run, single-year averages we compute in this study.
Our results align with existing research on the economic impacts of high temperatures [6, 25, 26] which found that productivity losses are the result of lower crop yields or impaired worker performance; we expand on this by showing that extreme heat also influences travel decisions. For example, anecdotal evidence suggests that farm workers avoid outdoor labor when temperatures exceed $40^{\circ}$ [49], and our data confirms this while pointing to broader changes in both urban and rural areas.
In India and Indonesia, respectively, we observe a noticeable decline in activity levels when average temperatures exceed 33C and 35C. We use rolling averages to show that, in India, while longer heat waves have stronger effects, activity resumes when high temperatures persist for a week or more. While extreme heat never shows a positive effect on mobility in our data, the negative effect converges to 0 when the temperature averages more than $40^{\circ}$ for 6 or 7 days. We are unable to examine the mechanisms by which this occurs but if it is the result of individuals needing to return to work, it adds to the importance of cooling efforts.
However, the effects we observe are not equally distributed. They are more pronounced in economically disadvantaged areas, indicating that these regions are more vulnerable to extreme weather disruptions. We find in India that foot traffic in areas characterized by high deprivation fell faster at higher temperatures than foot traffic in those with low deprivation. Changes in activity will result from combinations of forces: the ability of the wealthy to keep cool with air conditioning, for instance, might buoy activity in wealthy areas while pressures to earn might sustain activity in poor areas. Our findings suggest that for the poor, the expected gains from working are less than the dangers from working at high temperatures. This is especially concerning because residents in these areas are more likely to hold jobs that require physical labor, making adaptation more difficult.
Finally, by modeling future climate scenarios, we estimate how human activity will be affected in the years to come. These projections allow us to better understand the evolving situation, not just due to changing weather patterns but also due to broader shifts, such as the planned relocation of Indonesia’s capital from Java to the island of Borneo. Heat waves in India’s Rajasthan and Indonesia’s Sumatra will have the strongest effects on mobility, with temperatures exceeding $40^{\circ}$ more than 100 days per year in both, changes that are associated with 10% declines in activity according to our models.
A key open question remaining after our study is whether or not populations will gradually adapt with rising temperatures. This has relevance to our results because some places will have stronger heat waves, like Mexico’s arid North, and others will have higher but steadier temperatures, like its tropical South. Recent work on economic growth and public health looking across 5 decades suggests the answer is no: for most variables, responses to extreme heat have not changed [50]. An important finding in our work is that more deprived areas see stronger effects than less deprived ones, which has implications for adaptation: these areas will be amongst the last to adopt air conditioning. In the coming decades, our results may hold for vulnerable populations even if others are able to adapt.
Methods
We use origin-destination matrices for three nations—India, Indonesia and Mexico—spanning 2019. These data contain flows between tessellated units, called geohashes, which are on average $\sim 22\text{km}^{2}$. Aggregated data were provided by Spectus Social Impact as part of the Netmob 2024 conference. Data are collected with the informed consent of anonymous users who have opted in to anonymized data collection for research purposes. In order to further preserve the privacy of users, all data have been aggregated by the data provider to the geohash5 level and do not include any individual-level data records.
For each geohash5 in the data, using Google Earth Engine [51] we also gather 365 days of weather data from the ERA5 daily climate aggregates [42]. These data gives us temperature, precipitation and solar radiation per day per geohash5, and it allows us to observe differences within and between regions and does not rely on weather station data. We acquire modeled estimates of these variables for each day of 2050 using spatially disaggregated data from the Coupled Model Intercomparison Project [47], an international collaboration to produce estimates of Earth’s climate through 2100. These datasets both resolve to $\sim 775\text{km}^{2}$, and thus also capture spatial heterogeneity in weather conditions. Finally, we attach data—via zonal statistics—on population [45] and deprivation [46] to each geohash5, using gridded measures estimated from remote sensing.
Modeling the temperature gradient
We begin using a generalized additive model (GAM) to measure the partial effect of temperature at different levels [52]. We build a directed acyclic graph (DAG) to determine what controls are necessary in our model, shown in Fig. 2A. In addition to temperature, solar radiation and precipitation could also modulate activity, with rain dampening activity and sun heightening, so we control for these along with day-of-year. Day-of-year is modeled with a cyclic cubic spline, which captures seasonality by allowing for variation by day while forcing the value of the spline at the start to equal its value at the end. Because geography and area characteristics will also influence activity, we use geohash5 fixed effects, allowing the intercept to vary according to the unique activity profile of each geographic unit. Together, these spatial and temporal controls constitute the minimally sufficient adjustment set. While they are not confounds, adjusting for day-of-week and holiday effects improves the precision of our estimates. The resulting equation is
$T=\beta_{0}+f_{1}(x_{1})+f_{2}(x_{2})+\mathbf{X}\boldsymbol{\beta}+\varepsilon$ (1)
where $T$ is the number of trips ending in the geohash5, $f_{1}(x_{1})$ represents the smooth function of temperature, $f_{2}(x_{2})$ represents the smooth function of day (with a cyclic cubic spline), $\mathbf{X}$ is a matrix of control variables, and $\boldsymbol{\beta}$ is a vector of coefficients for the control variables. $\varepsilon$ is the error term.
Modeling the time series
Because GAMs do not have an explicit time component, we then model the time series as a robustness check. To analyze the relationship between activity and temperature while accounting for temporal dependencies, we employ an Autoregressive Integrated Moving Average model with Exogenous variables (ARIMAX) [53]. This model combines autoregressive terms, moving average components, and exogenous factors to capture the complex dynamics of the time series. Specifically, our model incorporates AR(1) and AR(7) terms to account for immediate changes and day-of-the-week patterns in activity. The moving average component allows the model to consider the impact of past shocks and manages seasonality. Temperature is included as an exogenous factor with both linear and quadratic terms to capture potential nonlinear effects on activity. Our model is specified as
$T_{t}=c+\phi_{1}T_{t-1}+\phi_{7}T_{t-7}+\beta_{1}\text{Temp}_{t}+\beta_{2}\text{Temp}_{t}^{2}+\theta_{1}\varepsilon_{t-1}+\varepsilon_{t}$ (2)
where $T_{t}$ represents economic activity measured as Trips at time $t$, $c$ is a constant term, $\phi_{1}$ and $\phi_{7}$ are autoregressive coefficients, $\beta_{1}$ and $\beta_{2}$ are coefficients for linear and quadratic temperature effects, $\text{Temp}_{t}$ is the temperature at time $t$, $\theta_{1}$ is the moving average coefficient, and $\varepsilon_{t}$ is the error term. This model structure allows us to simultaneously account for time-dependent patterns in economic activity and the influence of temperature, providing a comprehensive framework for analyzing the complex relationships within our data.
We compute the average coefficient for $\text{Temperature}^{2}$, which indicates the degree to which the effect of temperature curves at extremes, and we identify the ideal temperature by making predictions using the model for each time series at all temperatures and observing the point at which predictions begin to fall. We run a sensitivity analysis because not all time series span the full length of the year, showing the average coefficient at various minimum thresholds.
Projecting into the future
In order to understand how rising temperatures will impact behaviors, we take two approaches. First, we take the covariates from Eq. 1 but switch the temperatures on any given day for temperatures from the same day-of-year in 2050 using CMIP6 SSP 8.5 [47]. This gives us an aggregate estimate, incorporating both warmer winters (which might be good) and hotter summers (which might be bad). Second, we explore specific temperature thresholds that our modeling has shown to exert strong effects on mobility and explore how often these thresholds will be crossed according to CMIP6 SSP 8.5 estimates in 2050. To see estimates of temperatures in our 3 study countries through 2100, see Fig. S4.
Acknowledgements
The authors would like to thank the Spectus and the NetMob team for making the data available, as well as Wenlan Zhang for responding to questions and updating the data on multiple occasions.
Author contributions statement
A.R. Conceptualization, methodology, investigation, writing, reviewing, editing; O.H. Conceptualization, methodology, investigation, writing, reviewing, editing.
Data and code availability
Mobility data are available as part of the NetMob 2024 Data Challenge. CMIP6 projections are accessible via Google Earth Engine. All code is available upon request.
Competing interests
The authors declare no conflict of interest.